OpenStudy (anonymous):

What are the next three terms of the sequence 1, -3, 9, -27,… A. -54. 162, -486 B. -81, 243, -729 C. 81, -243, 729 D. 57, -162, 486

OpenStudy (solomonzelman):

\(\Large\color{blue}{ \rm ~1 \times (-3)=-3}\) \(\Large\color{blue}{ \rm ~(-3) \times (-3)=9}\) \(\Large\color{blue}{ \rm ~9 \times (-3)=-27}\) \(\Large\color{blue}{ \rm ~(-27) \times (-3)=81}\) \(\Large\color{blue}{ \rm ~81 \times (-3)=-243}\) \(\Large\color{blue}{ \rm ~(-243) \times (-3)=729}\)